### v4

@@ Abstract
We estimate the numbers of always takers, compliers, defiers, and never takers in the sample of people in an experiment, rather than a hypothetical population from which it was drawn, using structure from the randomization design. Our data include only a binary intervention and outcome. We develop a visualization to show that samples with defiers can sometimes generate the data in more ways than samples without defiers, yielding a higher design-based likelihood. We propose a maximum likelihood decision rule that can harness this evidence, which is not captured by standard hypothesis tests, and we provide optimality conditions. We illustrate the output of our decision rule for all possible data in samples of 50 and 200 with half in intervention, demonstrating a pattern in when the MLE includes defiers despite a positive average effect. We provide insights into effect heterogeneity in two published experiments with interventions that could plausibly backfire for some people despite statistically significant positive average effects on takeup of desirable health behaviors. In both, our 95% credible sets include the estimated Fréchet bounds, demonstrating that evidence is weak. Yet, our MLE includes no defiers in one; in the other, the MLE includes a count of defiers equal to the estimated upper Fréchet bound, over 18% of the sample. The MLE can support a monotonicity assumption or a specific alternative as a step toward improving the average effect of future interventions by targeting them away from noncompliers. Our dbmle package, compatible with Python and Stata, implements our statistics.
@@ 1 Introduction
We begin with a stylized example that demonstrates how a design-based model of an experiment can reveal evidence beyond the average effect. In Section 2, we formalize the model and present a likelihood function for the joint distribution of potential outcomes in the sample, which we use to construct a maximum likelihood decision rule. In Section 3, we apply our maximum likelihood decision rule to two real-world experiments. Section 4 discusses our contributions to the literature. Section 5 presents implications for econometric and applied research.
@@
Suppose you are considering an intervention to increase takeup of a desired health behavior. For example, the desired health behavior could be to quit smoking, and the intervention could be a payment. You are worried that your intervention could backfire, even if it increases takeup on average. Some people who would otherwise quit smoking could be dissuaded from doing so by a payment. You are also concerned that you would rather not waste payments on people who would quit smoking or continue smoking regardless.
A colleague tells you that you can use an experiment to estimate the average effect of the intervention in the population, but it cannot provide evidence that the intervention backfires if the average effect is positive. You are worried that even if the estimated average effect is positive, the intervention could backfire for some people and be ineffectual for others. Rather than focusing on the average effect in the population, you focus your attention on the distribution of effects within the fixed sample of subjects in an experiment.
You design an experiment with six people such that you randomly assign exactly three people to intervention and three people to control. You run the experiment, and you estimate the marginal distributions of outcomes in intervention and control: two people have taken up the behavior in intervention, and one person has taken up the behavior in control. These results seem promising because takeup is higher in intervention than control, yielding an average effect of 1/3 on a control takeup rate of 1/3.
@@
Why did you observe the data you observed? What would you have seen under the other possible randomized assignments? Can the experiment reveal any evidence about effect heterogeneity? In more technical terms, what is the joint distribution of “potential outcomes” (neyman1923; rubin1974; rubin1977) in the sample? Counterfactual questions like these have attracted recent interest in the study of causal inference (gelman2013; pearl2018; imbens2020; dawid2022).
@@
We assume, as in design-based inference, that the potential outcomes of each person are fixed, and therefore the joint distribution of potential outcomes in the sample is fixed, before randomization. Following angrist1996, we classify people into types based on their potential outcomes in intervention and control: “always takers” take up the desired health behavior in intervention and control (effect=0); “compliers” take up in intervention but not control (effect=1); “defiers” do not take up in intervention but take up in control (effect=-1); and “never takers” do not take up in intervention or control (effect=0). We specify the joint distribution of potential outcomes with the numbers of always takers, compliers, defiers, and never takers in the sample. The joint distribution of potential outcomes provides more information than the average effect; it provides the full distribution of effects and disambiguates between zero effects from always and never takers.
@@
What is your best estimate of the numbers of always takers, compliers, defiers, and never takers in the experiment you ran with six people? We estimate with our design-based maximum likelihood decision rule that the sample includes four compliers and two defiers, and we develop a novel visualization of potential outcomes to illustrate the logic behind our answer in Figure 1. We represent each of the six people in the sample with a circle. The left half of each circle represents the potential outcome in intervention, and the right half represents the potential outcome in control. Orange represents “takeup,” white represents “no takeup,” and grey represents “unobserved.”
@@
The data from your experiment, depicted in the box at the top of Figure 1, show takeup among two of three people in intervention and one of three people in control. Therefore, the estimated marginal distribution of the potential outcome in intervention is 2/3 takeup, and the estimated marginal distribution of the potential outcome in control is 1/3 takeup. These two marginal distributions define the “estimated Fréchet set.” We refer to all joint distributions of potential outcomes with the same marginal distributions as members of the same “Fréchet set.”
@@
The three rows of Figure 1 give the three joint distributions of potential outcomes with the same marginal distributions as the estimated Fréchet set: half circles in which the potential outcomes are unobserved in the data are filled in so that they have the same marginal distributions as the data. Four of the six circles have orange on the left (takeup occurs among 2/3 of people in intervention), and two of the six circles have orange on the right (takeup occurs among 1/3 of people in control). As shown in the three rows, there are three ways to combine the colored half circles into colored full circles, each yielding a different number of defiers (circles with white on the left and orange on the right). This depiction demonstrates how the estimated Fréchet set determines the estimated Fréchet bounds on defiers, which extend here from a lower bound of 0 to an upper bound of 2. How do you decide among these three possible estimates?
Figure 1: In an Experiment with Six People, the Likelihood Varies Among Joint Distributions of Potential Outcomes with the Same Marginal Distributions as the Data (2/3 Takeup in Intervention, 1/3 Takeup in Control) and is Maximized with Four Compliers and Two Defiers
@@
The joint distribution of potential outcomes in the first row is consistent with the imbens1994 assumption of “monotonicity” in the sample, whereby there can be compliers or defiers, but not both. Since the estimated effect of 1/3 is positive, you assume away defiers. You then deduce that the person who takes up in control must be an always taker and the person who does not take up in intervention must be a never taker. Assuming that the randomization guarantees that the count of each type is the same in each arm, the sample includes two always takers and two never takers. The remaining two people must be compliers, so the joint distribution of potential outcomes consistent with the monotonicity assumption includes two always takers, two compliers, and two never takers. However, you are uncomfortable assuming monotonicity because you are concerned that there could be defiers, and you would prefer to decide based on evidence.
@@
The joint distributions of potential outcomes in the next two rows include defiers. The one in the second row includes one always taker, three compliers, one defier, and one never taker. The one in the third row includes four compliers and two defiers. There is a rationale in the literature for deciding on the distribution in the second row—it has one defier, the midpoint (li2019) of the Fréchet bounds on defiers. It can therefore seem surprising that we decide on the distribution in the third row with two defiers.
@@
Each row in Figure 1 provides a rationale for our decision using a “design-based model” that combines the design of the randomization with a model of potential outcomes to specify a data generating process and use it to generate all possible data. As depicted in the top of each row, there are 20 possible randomized assignments of six people that result in three in intervention ($6\text{ choose }3=20$). Each of the six people in the experiment has a separate row. The person’s type and randomized assignment to intervention or control determines the data the experiment reveals, which can be counterfactual to what you observed. The assignments that generate the data you observed—2/3 takeup in intervention and 1/3 takeup in control—are in lighter shading.
@@
The number of randomized assignments that generate the data you observed gives the numerator of the design-based likelihood, reported in the third column. Paraphrasing the baby board book “Statistical Physics for Babies” (ferrie2017), which provides style inspiration for our visualization, physicists refer to the number of ways that you could have seen what you have seen as “entropy.” The design-based likelihood is equal to the entropy—the number of assignments that generate the data (shown in lighter shading)—divided by the number of possible assignments. The first row, with zero defiers, generates the data in eight of 20 possible assignments, so the likelihood is $8/20=40\%$. The second row, with one defier, generates the data in six of 20 possible assignments, so the likelihood is $6/20=30\%$. The last row, with two defiers, generates the data in 12 of 20 possible assignments, so the likelihood is $12/20=60\%$.
@@
Why do some joint distributions of potential outcomes generate the data in more randomized assignments than others and therefore have higher likelihoods than others? The bottom half of each row sorts the six people by type and separately shows how the always takers, compliers, defiers, and never takers must be randomized into intervention and control to generate the data. In the first row, there are three types—always takers, compliers, and never takers—and there must be a balance between intervention and control within each type to generate the data. In the second row, there are four types—always takers, compliers, defiers, and never takers—and there must be an imbalance between intervention and control within each type to generate the data. In the third row, there are only two types—compliers and defiers—and there must be a balance between intervention and control within each type to generate the data.
@@
Likelihoods are higher when the people who generate the data are as similar as possible in two senses: they belong to fewer types, and they have a counterpart of the same type in the opposite arm such that those of the same type are balanced between intervention and control. The people who generate the data are most similar in the third row: they belong to the fewest number of types (only two), and those of the same type are balanced between intervention and control. Comparison of the first and third rows illustrates how having fewer types can increase the likelihood. In the first row, there are two ways to randomize one of two always takers into intervention (2 $2\text{ choose }1$=2), two ways to randomize one of two compliers into intervention (2 $2\text{ choose }1$=2), and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are eight assignments that generate the data ($2\times 2\times 2=8$). In the third row, there are six ways to randomize two of four compliers into intervention ($4\text{ choose }2=6$) and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are 12 assignments that generate the data ($6\times 2=12$). Joint distributions of potential outcomes other than those in Figure 1 can generate the data (the true joint distribution of potential outcomes need not preserve the estimated marginal distributions), but by exhaustive grid search, you determine that the global maximizer of the likelihood is the one depicted in the third row, which only includes two types.
By our proposed maximum likelihood decision rule, you conclude that your best estimate for the joint distribution of potential outcomes in the sample consists of four compliers and two defiers. Using this distribution, you recover the estimated average effect of 1/3—of the six people in the experiment, the four compliers have an effect of positive one, and the two defiers have an effect of negative one ($(4\times 1+2\times-1)/6=1/3$). The design-based model of the experiment enables you to learn about not just the average effect of your intervention, but also about something more primitive: the joint distribution of potential outcomes, which yields the entire distribution of effects.
@@ 2 A Design-Based Model of an Experiment
2.1 A Design-Based Model Implies a Likelihood Function for the Joint Distribution of Potential Outcomes in the Sample
In this section, we formalize the design-based model of an experiment presented in our stylized example. A sample consists of $n$ subjects, each of whom has a binary potential outcome in intervention $Y_{I}\in\{0,1\}$ and a binary potential outcome in control $Y_{C}\in\{0,1\}$, where $1$ represents takeup and 0 represents no takeup. A subject’s observed outcome depends only on their own potential outcomes and their inclusion in the intervention or control arm, ruling out network-type effects through a “no interference” (cox1958) or “stable unit treatment value” (rubin1980) assumption.
Subjects are randomly assigned to intervention ($Z=I$) or control ($Z=C$), and assignment to intervention $Z$ is independent of a subject’s potential outcomes $(Y_{I},Y_{C})$. The observed outcome $Y$ is:
@@
Subjects in the sample belong to one of four types or “principal strata,” defined by combinations of potential outcomes (frangakis2002). Let $\theta_{11}$ represent the number of always takers $(Y_{I}=1,Y_{C}=1)$, $\theta_{10}$ the number of compliers $(Y_{I}=1,Y_{C}=0)$, $\theta_{01}$ the number of defiers $(Y_{I}=0,Y_{C}=1)$, and $\theta_{00}$ the number of never takers $(Y_{I}=0,Y_{C}=0)$. The collection of these four counts ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ constitutes the joint distribution of potential outcomes in the sample, which we write as counts rather than shares to ease the following likelihood notation. The experimental data ${\boldsymbol{X}}=(X_{I1},X_{I0},X_{C1},X_{C0})$ consist of the counts of subjects who take up in intervention $X_{I1}$, who do not take up in intervention $X_{I0}$, who take up in control $X_{C1}$, and who do not take up in control $X_{C0}$.
@@
We adopt a “design-based” approach, in which the joint distribution of potential outcomes $\boldsymbol{\theta}$} is fixed, but unknown, and all randomness in the experimental data $\boldsymbol{X}$} comes from the random assignment of subjects to either intervention or control. As shown in Appendix A, randomization implies a distribution for the experimental data $\boldsymbol{X}$} given the joint distribution of potential outcomes $\boldsymbol{\theta}$}, yielding a likelihood expression. For Bernoulli randomization, also known as simple randomization or independent and identically distributed (IID) randomization, with probability $p$ of assignment to intervention, the likelihood is:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}}$
$\qquad\qquad\times p^{x_{I1}+x_{I0}}(1-p)^{x_{C1}+x_{C0}},$
where the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ restricts $i$ such that the binomial coefficients remain well defined.
For a completely randomized experiment with $m$ subjects randomized into intervention, the likelihood is:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$.
While these randomization procedures imply different distributions for the experimental data, once we condition on the experimental data, the likelihoods are both proportional to:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (2)
This last expression appears in copas1973.
@@
This likelihood function coincides with the likelihoods we derive in the introductory example of Section 1. For the first distribution with zero defiers and the third distribution with two defiers in Figure 1, there is a single term in the likelihood summation. For the second distribution with one defier, there are two.
@@
Given a pair of marginal distributions of potential outcomes in intervention and control, there are numerous joint distributions consistent with them. boole1854, hoeffding1940, and frechet1957 derive bounds on the possible copulas connecting marginal distributions of random variables into joint distributions. By applying these bounds to a pair of marginal distributions of potential outcomes, we can determine the set of all joint distributions of potential outcomes consistent with the pair of marginal distributions. We refer to the set of joint distributions of potential outcomes consistent with a given pair of marginal distributions as the “Fréchet set” of those marginal distributions, $\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)$. In Figure 1, the three rows comprise the Fréchet set for $\theta_{1\bullet}=4$ and $\theta_{\bullet 1}=2$. We index the distributions in a Fréchet set by their number of defiers, $\theta_{01}$, which can take values in the range
@@
These inequalities provide the so-called “Fréchet bounds” on the number of defiers. Within these bounds, each value for the number of defiers implies a unique number of always takers, never takers, and compliers consistent with the given marginals:
$\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)=\Big\{{\boldsymbol{\theta}}\in\mathbb{Z}^{4}_{+}:\,$ $\theta_{01}\in[\max\{0,-(\theta_{1\bullet}-\theta_{\bullet 1})\},\min\{\theta_{\bullet 1},n-\theta_{1\bullet}\}],$
$\theta_{10}=\theta_{1\bullet}-\theta_{\bullet 1}+\theta_{01},$
$\theta_{11}=\theta_{\bullet 1}-\theta_{01},$
$\theta_{00}=n-\theta_{1\bullet}-\theta_{01}\Big\}.$
The bounds on the number of defiers $\theta_{01}$ ensure that these counts are non-negative and sum to $n$.
@@
We can form estimates of the marginal distributions of potential outcomes in intervention and control directly using the observed takeup rates in intervention and control:
$\hat{\theta}_{1\bullet}({\boldsymbol{x}})$ $=n\left(\frac{x_{I1}}{x_{I1}+x_{I0}}\right)$
$\hat{\theta}_{\bullet 1}({\boldsymbol{x}})$ $=n\left(\frac{x_{C1}}{x_{C1}+x_{C0}}\right).$ (4)
@@
When we report these estimates, we round them if necessary for comparison to our count-based distributions. In our design-based model of an experiment, these are unbiased estimators of the sample marginal distributions of potential outcomes for a completely randomized experiment (before rounding). In a standard sampling-based model of an experiment, dividing by $n$ yields consistent estimates of the population marginal distributions of potential outcomes.
@@
It is well-known that, in the sampling-based model of an experiment, the takeup counts in intervention and control contain no information about the joint distribution of potential outcomes in the population beyond their marginal distributions. In Appendix B, we recreate this result by showing that the likelihood of the population-level joint distribution of potential outcomes is constant, conditional on the population-level marginal distributions. In other words, the likelihood function is flat across every population-level Fréchet set.

### v5

@@ Abstract
Using only a binary intervention and outcome and the design of the randomization within an experiment, we construct a design-based likelihood of the joint distribution of potential outcomes in the sample—the numbers of always takers, compliers, defiers, and never takers. We develop a visualization to show that samples with defiers can sometimes generate the data in more ways than samples without, yielding a higher likelihood. This likelihood can vary within the Fréchet bounds, even though the traditional likelihood does not. Evidence is weak, but it exists, as we illustrate with health applications and our dbmle package.
@@ 1 Introduction
We begin with a stylized example that demonstrates how a design-based model of an experiment can reveal evidence beyond the average effect. In Section 2, we formalize the model and present a likelihood function for the joint distribution of potential outcomes in the sample, which we use to construct maximum likelihood estimates. In Section 3, we use our design-based likelihood in two real-world health applications. Section 4 discusses our contributions to the literature. Section 5 presents implications for econometric and applied research.
@@
You are a doctor at a rural health clinic, and you have six patients who are smokers. You would like them to quit smoking, so you consider offering them a payment contingent on quitting. You are optimistic that the payments would induce some patients to quit. However, your psychologist friend warns you that payments can crowd out intrinsic motivations (Gneezy and Rustichini, 2000). Your economist friend tells you that traditional economic models assume away the possibility (Björklund and Moffitt, 1987; Imbens and Angrist, 1994; Heckman and Vytlacil, 1999; Vytlacil, 2002). But you have taken an oath to “do no harm.” Even if the payment works on average, that is little comfort if some of your patients would actually be harmed. The more patients who would quit otherwise but would not because of your payment, the worse you would feel. For patients who would quit regardless, your concern is more modest—you would rather not waste the payments. Because the payment is contingent on quitting, you would not waste payments on any patients who would remain smoking regardless of the payment, but as their doctor, you would still feel bad that your intervention cannot reach them. You would really like to know the count of patients of each of the four types.
You know that experiments are a tool to reveal causal evidence on the average effect. You also know that with traditional inference, you can obtain Fréchet bounds (Boole, 1854; Hoeffding, 1940; Fréchet, 1957) on the distribution of effects in the population from which the smokers in your clinic were drawn. However, since your responsibilities are to your own patients, you do not focus on the population. You would like to estimate the full distribution of effects in your fixed sample of patients.
You plan an experiment, and you are deliberate about the design of your randomization. You choose a “completely randomized” design in which exactly three patients will be assigned to the payment intervention. You implement this design by using a computer to effectively pick three names out of a hat, without replacement.
You run the experiment, and the results are promising in terms of the average effect. Two of the three patients in intervention quit smoking; only one of the three patients in control quits. The intervention mean is 2/3, and the control mean is 1/3, so the average effect of the payment intervention is an increase in the quit rate of 1/3.
@@
Why did you observe the data you observed? What would you have seen under the other possible randomized assignments? Can the experiment reveal any evidence about effect heterogeneity? In more technical terms, what is the joint distribution of “potential outcomes” (Neyman, 1923; Rubin, 1974, 1977) in the sample? Counterfactual questions like these have attracted recent interest in the study of causal inference (Gelman and Imbens, 2013; Pearl and Mackenzie, 2018; Imbens, 2020; Dawid and Musio, 2022).
@@
We assume, as in design-based inference, that the potential outcomes of each person are fixed, and therefore the joint distribution of potential outcomes in the sample is fixed, before randomization. Following Angrist, Imbens, and Rubin (1996), we classify patients into types based on their potential outcomes in intervention and control: “always takers” take up the desired health behavior—they quit smoking—in intervention and control (effect=0); “compliers” take up in intervention but not control (effect=1); “defiers” do not take up in intervention but take up in control (effect=-1); and “never takers” do not take up in intervention or control (effect=0). We specify the joint distribution of potential outcomes with the numbers of always takers, compliers, defiers, and never takers in the sample. The joint distribution of potential outcomes provides more information than the average effect; it provides the full distribution of effects and disambiguates the zero effects of always takers from those of never takers.
@@
We demonstrate that a design-based model of your experiment can reveal evidence on the numbers of always takers, compliers, defiers, and never takers in your clinic, beyond the evidence revealed by the estimated average effect and the resulting Fréchet bounds. We develop a novel visualization of potential outcomes to illustrate how in Figure 1. We represent each of the six patients in the sample with a circle. The left half of each circle represents the potential outcome in the payment intervention, and the right half represents the potential outcome in the no payment control. Orange represents “takeup” of the desired health behavior—quitting smoking; white represents “no takeup,” and grey represents “unobserved.”
Figure 1: In an Experiment with Six Patients, the Likelihood Varies Among Joint Distributions of Potential Outcomes with the Same Marginal Distributions as the Data (2/3 Takeup in Intervention, 1/3 Takeup in Control) and is Maximized with Four Compliers and Two Defiers
@@
The data from your experiment, depicted in the box at the top of Figure 1, show takeup among two of three patients in intervention and one of three patients in control. Therefore, the estimated marginal distribution of the potential outcome in intervention is 2/3 takeup, and the estimated marginal distribution of the potential outcome in control is 1/3 takeup. These two marginal distributions define the “estimated Fréchet set.” We refer to all joint distributions of potential outcomes with the same marginal distributions as members of the same “Fréchet set.”
@@
The three rows of Figure 1 give the three joint distributions of potential outcomes with the same marginal distributions as the estimated Fréchet set: half circles in which the potential outcomes are unobserved in the data are filled in so that they have the same marginal distributions as the data. Four of the six circles have orange on the left (takeup occurs among 2/3 of patients in intervention), and two of the six circles have orange on the right (takeup occurs among 1/3 of patients in control). As shown in the three rows, there are three ways to combine the colored half circles into colored full circles, each yielding a different number of defiers (circles with white on the left and orange on the right). This depiction demonstrates how the estimated Fréchet set determines the estimated Fréchet bounds on defiers, which extend here from a lower bound of 0 to an upper bound of 2. Is there any variation in the design-based likelihood across the joint distributions of potential outcomes in the three rows?
@@
The joint distribution of potential outcomes in the first row is consistent with the Imbens and Angrist (1994) assumption of “monotonicity” in the sample, whereby there can be compliers or defiers, but not both. Since the estimated effect of 1/3 is positive, you assume away defiers. You then deduce that the person who takes up in control must be an always taker and the person who does not take up in intervention must be a never taker. Assuming that the count of each type is the same in each arm, the sample includes two always takers and two never takers. The remaining two patients must be compliers, so the joint distribution of potential outcomes consistent with the monotonicity assumption includes two always takers, two compliers, and two never takers. However, you are uncomfortable assuming monotonicity because you are concerned that there could be defiers, and you would prefer to rely on evidence.
@@
The joint distributions of potential outcomes in the next two rows include defiers. The one in the second row includes one always taker, three compliers, one defier, and one never taker. The one in the third row includes four compliers and two defiers. There is a rationale in the literature for assuming the distribution in the second row—it has one defier, the midpoint (Li and Pearl, 2019) of the Fréchet bounds on defiers. It can therefore seem surprising that evidence yields a rationale for assuming the distribution in the third row with two defiers.
@@
Each row in Figure 1 depicts evidence from a “design-based model” that combines the design of the randomization with a model of potential outcomes to specify a data generating process and use it to generate all possible data. As depicted in the top of each row, there are 20 possible randomized assignments of six patients that result in three in intervention ($6\text{ choose }3=20$). By the design of your experiment, each of these possible randomized assignments has the same probability, so you can make the assumption that they each have the same probability very compelling through careful implementation.
Within each main row, each of the six patients in the experiment has a separate row. The person’s type and randomized assignment to intervention or control determines the data the experiment reveals, which can be counterfactual to what you observed. The assignments that generate the data you observed—2/3 takeup in intervention and 1/3 takeup in control—are in lighter shading.
@@
The number of randomized assignments that generate the data you observed gives the numerator of the design-based likelihood of the joint distribution of potential outcomes, reported in the third column. Paraphrasing the baby board book “Statistical Physics for Babies” (Ferrie, 2017), which provides style inspiration for our visualization, physicists refer to the number of ways that you could have seen what you have seen as “entropy.” The design-based likelihood is equal to the entropy—the number of assignments that generate the data (shown in lighter shading)—divided by the number of possible assignments. The first row, with zero defiers, generates the data in eight of 20 possible assignments, so the likelihood is $8/20=40\%$. The second row, with one defier, generates the data in six of 20 possible assignments, so the likelihood is $6/20=30\%$. The last row, with two defiers, generates the data in 12 of 20 possible assignments, so the likelihood is $12/20=60\%$.
@@
Why do some joint distributions of potential outcomes generate the data in more randomized assignments than others and therefore have higher likelihoods than others? The bottom half of each row sorts the six patients by type and separately shows how the always takers, compliers, defiers, and never takers must be randomized into intervention and control to generate the data. In the first row, there are three types—always takers, compliers, and never takers—and there must be a balance between intervention and control within each type to generate the data. In the second row, there are four types—always takers, compliers, defiers, and never takers—and there must be an imbalance between intervention and control within each type to generate the data. In the third row, there are only two types—compliers and defiers—and there must be a balance between intervention and control within each type to generate the data.
@@
Likelihoods are higher when the patients who generate the data are as similar as possible in two senses: they belong to fewer types, and they have a counterpart of the same type in the opposite arm such that those of the same type are balanced between intervention and control. The patients who generate the data are most similar in the third row: they belong to the fewest number of types (only two), and those of the same type are balanced between intervention and control. Comparison of the first and third rows illustrates how having fewer types can increase the likelihood. In the first row, there are two ways to randomize one of two always takers into intervention (2 $2\text{ choose }1$=2), two ways to randomize one of two compliers into intervention (2 $2\text{ choose }1$=2), and two ways to randomize one of two never takers into intervention (2 $2\text{ choose }1$=2), so there are eight assignments that generate the data ($2\times 2\times 2=8$). In the third row, there are six ways to randomize two of four compliers into intervention ($4\text{ choose }2=6$) and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are 12 assignments that generate the data ($6\times 2=12$). Joint distributions of potential outcomes other than those in Figure 1 can generate the data; the true joint distribution of potential outcomes need not preserve the estimated marginal distributions. For example, the likelihood that the true average effect is zero and your sample contains only always and never takers is $9/20=45\%$. However, by exhaustive grid search, you determine that the global maximizer of the likelihood is the one depicted in the third row, which contains four compliers and two defiers.
As a doctor considering the distribution with the maximum likelihood, you are thrilled that it implies that the payment intervention would induce four of your six patients to quit smoking, which is even more promising than what you would conclude based on the average effect and a monotonicity assumption. You are also encouraged that it implies that you do not waste any payments on patients who would have quit regardless and that there are no patients who would continue smoking regardless. However, you are concerned that it implies that the payment would induce two patients who would otherwise have quit to continue to smoke.
@@ 2 A Design-Based Model of an Experiment
2.1 A Design-Based Model Implies a Likelihood for the Joint Distribution of Potential Outcomes in the Sample
Here, we formalize the design-based model of an experiment presented in our stylized example. A sample consists of $n$ subjects, each of whom has a binary potential outcome in intervention $Y_{I}\in\{0,1\}$ and a binary potential outcome in control $Y_{C}\in\{0,1\}$, where $1$ represents takeup and 0 represents no takeup. A subject’s observed outcome depends only on their own potential outcomes and their inclusion in the intervention or control arm, ruling out network-type effects through a “no interference” (Cox, 1958) or “stable unit treatment value” (Rubin, 1980) assumption.
Subjects receive a random assignment $Z$ to intervention ($Z=I$) or control ($Z=C$), which is independent of a subject’s potential outcomes $(Y_{I},Y_{C})$. The observed outcome $Y$ is:
@@
Subjects in the sample belong to one of four types or “principal strata,” defined by combinations of potential outcomes (Frangakis and Rubin, 2002). Let $\theta_{11}$ represent the number of always takers $(Y_{I}=1,Y_{C}=1)$, $\theta_{10}$ the number of compliers $(Y_{I}=1,Y_{C}=0)$, $\theta_{01}$ the number of defiers $(Y_{I}=0,Y_{C}=1)$, and $\theta_{00}$ the number of never takers $(Y_{I}=0,Y_{C}=0)$. The collection of these four counts ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ constitutes the joint distribution of potential outcomes in the sample, which we write as counts rather than shares to ease the following likelihood notation. The experimental data ${\boldsymbol{X}}=(X_{I1},X_{I0},X_{C1},X_{C0})$ consist of the counts of subjects who take up in intervention $X_{I1}$, who do not take up in intervention $X_{I0}$, who take up in control $X_{C1}$, and who do not take up in control $X_{C0}$.
@@
We adopt a “design-based” approach, in which the joint distribution of potential outcomes $\boldsymbol{\theta}$} is fixed, but unknown, and all randomness in the experimental data $\boldsymbol{X}$} comes from the random assignment of subjects to either intervention or control. As shown in Appendix A.1, randomization implies a distribution for the experimental data $\boldsymbol{X}$} given the joint distribution of potential outcomes $\boldsymbol{\theta}$}, yielding a likelihood expression. For an experiment with Bernoulli randomization (also known as simple randomization or independent and identically distributed randomization) or a completely randomized experiment, the likelihood is proportional to:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{j\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{j}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-j}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-j}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}}$, (2)
@@
where the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ restricts $j$ such that the binomial coefficients remain well defined. This expression appears in Copas (1973). The likelihoods we derive in the stylized example of Section 1 are also proportional to (2). For the first distribution with zero defiers and the third distribution with two defiers in Figure 1, there is a single term in the likelihood summation. For the second distribution with one defier, there are two.
@@
Given a pair of marginal distributions of potential outcomes in intervention and control, there are numerous joint distributions consistent with them. Boole (1854), Hoeffding (1940), and Fréchet (1957) derive bounds on the possible copulas connecting marginal distributions of random variables into joint distributions. By applying these bounds to a pair of marginal distributions of potential outcomes, we can determine the set of all joint distributions of potential outcomes consistent with the pair of marginal distributions. We refer to the set of joint distributions of potential outcomes consistent with a given pair of marginal distributions as the “Fréchet set” $\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)$ of those marginal distributions:
$\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)=\bigg\{{\boldsymbol{\theta}}\in\mathbb{Z}^{4}_{+}:\,$ $\theta_{11}+\theta_{10}=\theta_{1\bullet},\ \theta_{11}+\theta_{01}=\theta_{\bullet 1},\ \sum_{i,j\in\{0,1\}}\theta_{ij}=n\bigg\}.$
In Figure 1, the three rows comprise the Fréchet set for $\theta_{1\bullet}=4$ and $\theta_{\bullet 1}=2$. We index the distributions in a Fréchet set by their number of defiers, $\theta_{01}$, which can take values in the range
@@
These inequalities provide the so-called “Fréchet bounds” on the number of defiers.
@@
We can form estimates of the marginal distributions of potential outcomes in intervention and control using the observed takeup rates in intervention and control:
$\hat{\theta}_{1\bullet}({\boldsymbol{x}})=n\left(\frac{x_{I1}}{x_{I1}+x_{I0}}\right),\qquad\hat{\theta}_{\bullet 1}({\boldsymbol{x}})=n\left(\frac{x_{C1}}{x_{C1}+x_{C0}}\right).$ (4)
@@
In our design-based model of an experiment, these are unbiased estimators of the sample marginal distributions of potential outcomes for a completely randomized experiment. In a standard sampling-based model of an experiment, dividing by $n$ yields consistent estimates of the population marginal distributions of potential outcomes.
@@
It is well-known that, in the sampling-based model of an experiment, the takeup counts in intervention and control contain no information about the joint distribution of potential outcomes in the population beyond its marginal distributions. In Appendix A.2, we recreate this result by showing that the likelihood of the population-level joint distribution of potential outcomes is constant, conditional on the population-level marginal distributions. In other words, the likelihood function is flat within every population-level Fréchet set.